Let V be the set R+ of positive real number and define V1 "\bigoplus" V2=V1V2 and "\delta" "\bigodot" V1=V1"\delta" for all V1V2"\epsilon" V and "\delta" "\epsilon" R. Then show that v is a vector space over R


1
Expert's answer
2021-10-29T03:02:13-0400

A vector space over "\R" consist of a set "V" on which is defined of addition associated to elements "v_1\>and\>v_2\>of\>V,\>" "an \> element \>u\>and\>v\>of \>V"


And an operation of multiplication by scalars,associated to each element "\delta" of "\R" and to each element "v\>of\>V"

an element "\delta\>v\>of\>V"


"V" satisfies the following axioms

"1.\>v_1+v_2=v_2+v_1"

"2.\>(v_1+v_2)+v_3=v_1+(v_2+v_3)"

"3." There exists a 0 element"\>0\>of\>V" such that"\>v+0=v"

"4." Given any element "v\>of\>V" there exists "\>-v\>of\>V" with the property that

"v+(-v)=0"

"5." "(\delta _1+\delta_2)v=\delta_1v+\delta_2v"

"6. \>\delta(v_1+v_2)=\delta\>v_1+\delta\>v_2"

"7.\>\delta_1(\delta_2v)=(\delta_1\delta_2)v"

"8.\>Iv=v"

Where "I" is the multiplicative identity element of "\R"



Therefore "V" is a vector space over "\R"